# Pathwise Test-Time Correction for Autoregressive Long Video Generation

### v1

@@ Abstract
Distilled autoregressive diffusion models facilitate real-time short video synthesis but suffer from severe error accumulation during long-sequence generation. While existing Test-Time Optimization (TTO) methods prove effective for images or short clips, we identify that they fail to mitigate drift in extended sequences due to unstable reward landscapes and the hypersensitivity of distilled parameters. To overcome these limitations, we introduce Test-Time Correction (TTC), a training-free alternative. Specifically, TTC utilizes the initial frame as a stable reference anchor to calibrate intermediate stochastic states along the sampling trajectory. Extensive experiments demonstrate that our method seamlessly integrates with various distilled models, extending generation lengths with negligible overhead while matching the quality of resource-intensive training-based methods on 30-second benchmarks.
Machine Learning, ICML
@@ 1 Introduction
Video generation (Lu et al., 2025; Yesiltepe et al., 2025; Hong et al., 2023; Jia et al., 2025b) has advanced rapidly with the development of diffusion-based generative models (Kong et al., 2024; Wan et al., 2025; Hong et al., 2023; Ma et al., 2025a; Peebles and Xie, 2023; Rombach et al., 2022), which now enable the high-quality synthesis of complex motion (Zhu et al., 2024; Hu, 2024) and visual appearance (Guo et al., 2024; Zhang et al., 2025a). However, scaling these diffusion priors to extended video sequences remains a formidable challenge. Beyond the escalating computational costs associated with longer contexts, maintaining temporal coherence over extended horizons is difficult without incurring excessive latency, thereby limiting their deployment in real-time applications.
To overcome these limitations, recent studies (Yin et al., 2025a; Huang et al., 2025c) have shifted from bidirectional modeling to step-distilled autoregressive generation, enabling true real-time video synthesis. However, these methods remain constrained by cascading error accumulation: since each frame is conditioned on prior outputs, initial inaccuracies compound over time, resulting in temporal drift and long-horizon degradation. While recent extensions (Yi et al., 2025; Cui et al., 2026) like Rolling Forcing (Liu et al., 2025b), LongLive (Yang et al., 2025), Self-Forcing++ (Cui et al., 2025) , and WorldPlay (Sun et al., 2025) have achieved minute-level consistency through sink mechanisms and windowed DMD retraining, they necessitate substantial computational overhead for model fine-tuning. Consequently, a pivotal question arises: Can we improve the stability of autoregressive video generation purely at inference time, bypassing the need for retraining the base model?
Test-Time Optimization (TTO) (Wang et al., 2025; Sun et al., 2020) has emerged as a compelling alternative for enhancing video quality without the need for retraining. However, while effective for short-video synthesis (Yu et al., 2025b; Eyring et al., 2025), our toy experiments reveal that scaling TTO to long-horizon autoregressive generation faces a dual bottleneck consisting of the inherent difficulty in defining reward functions for long-range consistency and the extreme optimization sensitivity of distilled models. We observe that in these distilled models, even infinitesimal test-time gradients often trigger reward collapse and fail to mitigate cumulative error. Therefore, we propose Test-Time Correction (TTC), which is a training-free framework that shifts the paradigm from parameter-space optimization to sampling-space stochastic intervention. TTC is grounded in the insight that few-step distilled samplers are inherently stochastic as they perturb intermediate states with injected noise. This property implies that intermediate predictions are not fixed outcomes but rather malleable latent states that can be rectified by subsequent diffusion steps to align with the global initial context while preserving the underlying sampling distribution.
@@
Figure 2: Comparison of sampling strategies. The Original Path suffers from error accumulation, while the Sink-based Path collapses into a Sink Point (dynamic collapse). In contrast, our TTC strategy avoids these failures by employing reference-conditioned denoising and explicit Re-noising, effectively steering the trajectory away from the sink to preserve target distribution.
@@
Specifically, as shown in Figure 2, TTC applies a small number of correction steps along the stochastic sampling path, only after the global structure has stabilized. This delay prevents the generation from falling into sink-collapse (Cui et al., 2026), a phenomenon where newly generated frames repeatedly regress toward the sink frames instead of evolving naturally. At these chosen steps in the sampling path, TTC performs reference-conditioned denoising by utilizing the initial frame context to anchor a corrected clean prediction. Then, this corrected state is re-noised back to the variance level corresponding to the current timestep, which ensures that the intervention remains compatible with the expected noise distribution. By integrating correction into the stochastic sampling path of the autoregressive diffusion process rather than directly replacing the denoised prediction, this mechanism suppresses long-term error accumulation and temporal drift without retraining, while preserving high-fidelity temporal coherence over extended durations.
@@
In this work, we show that long-horizon stability in autoregressive video generation can be achieved through test-time intervention alone. Our method suppresses error accumulation with negligible computational overhead, without requiring any retraining. As a result, it extends the stable generation length of distilled autoregressive models from a few seconds to over 30 seconds, while achieving visual quality comparable to state-of-the-art training-based methods. Consistent improvements across multiple model architectures demonstrate that TTC is a robust and general solution for stabilizing distilled autoregressive diffusion models.
@@ 2 Related Work
Bidirectional Models for Video Generation. Diffusion models have been widely adopted for video generation, with recent works typically formulating video synthesis as a sequence-level joint denoising problem. Under this bidirectional diffusion paradigm, all frames are denoised simultaneously via spatiotemporal attention, enabling the model to leverage global temporal context and produce temporally coherent, high-fidelity videos (Blattmann et al., 2023; Yin et al., 2023; Jia et al., 2025a; Ma et al., 2025b; Zhang et al., 2025c; Huang et al., 2025b). Large-scale systems such as Hunyuan (Kong et al., 2024) and Wan (Wan et al., 2025) further demonstrate the effectiveness of this joint denoising formulation at scale. However, because the entire sequence must be processed as a whole during inference, this paradigm inherently precludes streaming or incremental generation, limiting its applicability in real-time and interactive scenarios.
@@
Autoregressive Models for Video Generation. Autoregressive Models for Video Generation. Autoregressive video diffusion models generate videos sequentially in a strict causal manner, conditioning each new frame or segment on the historical context of previously generated content (Chen et al., 2024; Yin et al., 2025a; Huang et al., 2025c; Liu et al., 2025b; Yang et al., 2025; Chen et al., 2025a; Ji et al., 2025; Teng et al., 2025; Deng et al., 2025; Guo et al., 2025). While this formulation naturally supports streaming generation with low initial latency, it is inherently susceptible to error accumulation, where minor deviations propagate and amplify across steps, leading to severe temporal drift and degraded coherence in long videos. To mitigate this issue, recent works introduce planning methods (Zhang et al., 2025b; Xiang et al., 2025) or explicit memory mechanisms (Sun et al., 2025; Chen et al., 2025b; Yu et al., 2025a; Cai et al., 2025; Huang et al., 2025a; HunyuanWorld, 2025), strategies that typically necessitate complex architectural modifications and extensive re-training.
Test-time Image/Video Generation. Test-time generation methods (Wang et al., 2025; Sun et al., 2020; Liang et al., 2025) aim to enhance the performance of pre-trained models directly during the inference phase. Test-time scaling improves quality by iteratively searching over multiple candidates, as seen in Video-T1 (Liu et al., 2025a) and EvoSearch (He et al., 2025), yet this comes at the price of prohibitive computational cost. Similarly, Test-time optimization refines generation via auxiliary parameter updates that necessitate instance-specific training, such as HyperNoise (Eyring et al., 2025), AutoRefiner (Yu et al., 2025b), and SLOWFAST-VGEN (Hong et al., 2025). In contrast, our approach distinguishes itself from both paradigms by being fully training-free, avoiding both the overhead of candidate search and the complexity of parameter optimization.
@@ 3.1 Background: Few-step Distilled Sampling
Under this formulation, each conditional distribution $p_{\theta}(x_{t}\mid S_{t})$ is no longer realized as a single deterministic mapping, but through a stochastic diffusion sampling trajectory defined over a sparse set of diffusion steps $\{T_{0}=0,T_{1},\dots,T_{K}=T_{\max}\}$. Specially, distilled video generation begins from Gaussian noise $x_{t}^{T_{\max}}\sim\mathcal{N}(0,I)$ and evolves progressively along this trajectory via a sequence of denoise–renoise transitions. Specifically, at each denoising step $T_{j}$, the generation process starts from a noisy latent state and applies the denoising network to produce an estimate of the underlying clean latent representation,
@@
Figure 4: Comparison of two toy test-time optimization variants based on LoRA fine-tuning.
3.2 Toy Experiment: Apply Test-time Optimization to Long Video Generation
Existing test-time optimization (TTO) methods improve generation quality by aligning the model distribution with a predefined reward function. Given a pre-trained generative model with output distribution $p^{\mathrm{base}}$, TTO typically defines a reward-weighted target distribution
@@ 3.2 Toy Experiment: Apply Test-time Optimization to Long Video Generation
To assess this idea, we conduct two toy experiments using direct LoRA fine-tuning (Hu et al., 2022) at test time, following HyperNoise (Eyring et al., 2025) and AutoRefiner (Yu et al., 2025b). Both variants use the same backbone and identical LoRA adapters, and differ only in their optimization objectives. The first variant fine-tunes LoRA with a standard denoising reconstruction loss on early frames across noise levels. The second variant replaces pixel-level reconstruction with a semantic consistency objective, enforcing similarity to early frames in pretrained feature spaces (Radford et al., 2021; Oquab et al., 2024).
@@ 4.1 Correctability along the Stochastic Sampling Path
Motivated by planning-based video generation models (Zhang et al., 2025b; Xiang et al., 2025), which relax strictly unidirectional prediction via cross-frame context, we consider applying test-time intervention at a single sampling step after structural stabilization. Specifically, at a designated step $j^{\star}$, we restrict the visible context state $S_{t}$ to include only the earliest frame, forcing the model to rely exclusively on the initial frame for subsequent appearance refinement and texture generation. This single-point correction process can be formalized as:
@@
4.2 Path-wise Test-time Correction
@@
Algorithm 1 Path-wise Test-time Correction
1:Input: Noise schedule $\{T_{J}>\cdots>T_{0}=0\}$; Generator $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}G_{\theta}$}; Evolving context $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{t}$}; Ref context $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{0}$}; Correction indices $\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\mathcal{J}^{\star}$}; Diffusion forward process $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\Psi$}
@@
3: Sample $x_{t}^{T_{J}}\sim\mathcal{N}(0,I)$ # Initial Noise
4:for $j=J$ down to 22 do
5:  $x_{t,0}^{T_{j}}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}G_{\theta}}\!\left(x_{t}^{T_{j}};\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{t}},\,j\right)$ # Initial prediction with $S_{t}$
@@
9:   $x_{t}^{T_{j-1},c}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\Psi}\!\left(x_{t,0}^{T_{j}},\,\epsilon_{t}^{T_{j-1}},\,T_{j-1}\right)$
10:   $x_{t,0}^{T_{j-1},c}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}G_{\theta}}\!\left(x_{t}^{T_{j-1},c};\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{0}},\,j-1\right)$ # Correct trajectory using $S_{0}$
@@
13:   $x_{t}^{T_{j-1}}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\Psi}\!\left(x_{t,0}^{T_{j-1},c},\,\tilde{\epsilon}_{t}^{T_{j-1}},\,T_{j-1}\right)$ # Inject new noise
14:   $x_{t,0}^{T_{j-1}}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}G_{\theta}}\!\left(x_{t}^{T_{j-1}};\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{t}},\,j-1\right)$ # Finalize step with $S_{t}$
@@
16:   $x_{t}^{T_{j-1}}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\Psi}\!\left(x_{t,0}^{T_{j}},\,\epsilon_{t}^{T_{j-1}},\,T_{j-1}\right)$
17:   $x_{t,0}^{T_{j-1}}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}G_{\theta}}\!\left(x_{t}^{T_{j-1}};\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{t}},\,j-1\right)$
@@
20:return $x_{t,0}^{T_{1}}$
Table 1: Comprehensive comparison with SOTA methods on prompt-conditioned 30-second video generation. We report Throughput (fps), VBench metrics, Color-shift metrics (L1, Correlation), and JEPA consistency (Standard Deviation, Difference).
Method Training Free Speed VBench Metrics Color-shift JEPA Consistency
Total fps Subject Consistency Background Consistency Dynamic Degree Motion Smoothness Imaging Quality Aesthetic Quality L1 $\downarrow$ Correlation $\uparrow$ Standard Deviation $\downarrow$ Difference $\downarrow$
Rolling Forcing 15.38 95.8 95.1 35.9 98.9 72.5 63.6 0.436 0.858 0.0162 0.201
LongLive - 95.5 95.4 44.5 98.8 71.7 65.0 0.701 0.724 0.0151 0.101
CausVid 15.79 91.2 91.4 50.8 98.1 70.2 63.5 1.047 0.451 0.0199 0.313
CV + Ours 10.53 93.2 93.3 69.5 97.6 70.1 63.5 0.607 0.778 0.0157 0.164
Self-Forcing 15.79 92.5 93.2 62.5 98.0 72.5 63.4 1.028 0.479 0.0145 0.191
SF + Ours 10.53 94.0 94.2 60.2 98.3 72.7 63.8 0.644 0.710 0.0108 0.170
Table 2: Comprehensive comparison with test-time scaling methods on prompt-conditioned 30-second video generation. We report Throughput (fps) and VBench metrics.
Method Train. Free Speed VBench Metrics
Total fps Sub. Cons. Bg. Cons. Dyn. Deg. Mot. Sm. Img. Qual. Aes. Qual.

### v2

@@ Abstract
Distilled autoregressive diffusion models facilitate real-time short video synthesis but suffer from severe error accumulation during long-sequence generation. While existing Test-Time Optimization (TTO) methods prove effective for images or short clips, we identify that they fail to mitigate drift in extended sequences due to unstable reward landscapes and the hypersensitivity of distilled parameters. To overcome these limitations, we introduce Test-Time Correction (TTC), a training-free alternative. Specifically, TTC utilizes the initial frame as a stable reference anchor to calibrate intermediate stochastic states along the sampling trajectory. Extensive experiments demonstrate that our method seamlessly integrates with various distilled models, extending generation lengths with a slight overhead while matching the quality of resource-intensive training-based methods on 30-second benchmarks.
Figure 1: 30-second video generation examples. Our method reduces error accumulation in CausVid and Self-Forcing, enabling longer and more stable videos with improved visual consistency. All samples are generated with the same random seed for fair comparison.
@@ 1 Introduction
Video generation [rewardforcing, infinityrope, Cogvideo, moga1] has advanced rapidly with the development of diffusion-based generative models [kong2024hunyuanvideo, wan2025, Cogvideo, Latte, DiT, LDM], which now enable the high-quality synthesis of complex motion [champ, Animateanyone] and visual appearance [animatediff, Proteus]. However, scaling these diffusion priors to extended video sequences remains a formidable challenge. Beyond the escalating computational costs associated with longer contexts, maintaining temporal coherence over extended horizons is difficult without incurring excessive latency, thereby limiting their deployment in real-time applications.
@@
Figure 2: Comparison of sampling strategies. The Original Path suffers from error accumulation, while the Sink-based Path collapses into a Sink Point. Our TTC strategy avoids this via reference-conditioned denoising and explicit Re-noising.
To overcome these limitations, recent studies [yin2025slow, huang2025self] have shifted from bidirectional modeling to step-distilled autoregressive generation, enabling true real-time video synthesis. However, these methods remain constrained by cascading error accumulation: since each frame is conditioned on prior outputs, initial inaccuracies compound over time, resulting in temporal drift and long-horizon degradation. While recent extensions [yi2025deep, lol] like Rolling Forcing [liu2025rolling], LongLive [yang2025longlive], Self-Forcing++ [Self-Forcing++] , and WorldPlay [sun2025worldplay] have achieved minute-level consistency through sink mechanisms and windowed DMD retraining, they necessitate substantial computational overhead for model fine-tuning. Consequently, a pivotal question arises: Can we improve the stability of autoregressive video generation purely at inference time, bypassing the need for retraining the base model?
TTO [wang2025test, sun2020test] has emerged as a compelling alternative for enhancing video quality without the need for retraining. However, while effective for short-video synthesis [yu2025autorefiner, eyring2025noise], our toy experiments reveal that scaling TTO to long-horizon autoregressive generation faces a dual bottleneck consisting of the inherent difficulty in defining reward functions for long-range consistency and the extreme optimization sensitivity of distilled models. We observe that in these distilled models, even infinitesimal test-time gradients often trigger reward collapse and fail to mitigate cumulative error. Therefore, we propose TTC, which is a training-free framework that shifts the paradigm from parameter-space optimization to sampling-space stochastic intervention. TTC is grounded in the insight that few-step distilled samplers are inherently stochastic as they perturb intermediate states with injected noise. This property implies that intermediate predictions are not fixed outcomes but rather malleable latent states that can be rectified by subsequent diffusion steps to align with the global initial context while preserving the underlying sampling distribution.
Specifically, as shown in Figure 2, TTC applies a small number of correction steps along the stochastic sampling path, only after the global structure has stabilized. This delay prevents the generation from falling into sink-collapse [lol], a phenomenon where newly generated frames repeatedly regress toward the sink frames instead of evolving naturally.
@@
At chosen steps in the sampling path, TTC performs reference-conditioned denoising by utilizing the initial frame context to anchor a corrected clean prediction. Then, this corrected state is re-noised back to the variance level corresponding to the current timestep, which ensures that the intervention remains compatible with the expected noise distribution. By integrating correction into the stochastic sampling path of the autoregressive diffusion process rather than directly replacing the denoised prediction, this mechanism suppresses long-term error accumulation and temporal drift without retraining, while preserving high-fidelity temporal coherence over extended durations.
@@
In this work, we show that long-horizon stability in autoregressive video generation can be achieved through test-time intervention alone. Our method suppresses error accumulation with a slight computational overhead, without requiring any retraining. As a result, it extends the stable generation length of distilled autoregressive models from a few seconds to over 30 seconds, while achieving visual quality comparable to state-of-the-art training-based methods. Consistent improvements across multiple model architectures demonstrate that TTC is a robust and general solution for stabilizing distilled autoregressive diffusion models.
@@ 2 Related Work
Bidirectional Models for Video Generation. Diffusion models have been widely adopted for video generation, with recent works typically formulating video synthesis as a sequence-level joint denoising problem. Under this bidirectional diffusion paradigm, all frames are denoised simultaneously via spatiotemporal attention, enabling the model to leverage global temporal context and produce temporally coherent, high-fidelity videos [blattmann2023align, yin2023nuwa, jia2025moga, ma2025tempomaster, zhang2025fast, Voyager]. Large-scale systems such as Hunyuan [kong2024hunyuanvideo] and Wan [wan2025] further demonstrate the effectiveness of this joint denoising formulation at scale. However, because the entire sequence must be processed as a whole during inference, this paradigm inherently precludes streaming or incremental generation, limiting its applicability in real-time and interactive scenarios.
@@
Autoregressive Models for Video Generation. Autoregressive video diffusion models generate videos sequentially in a strict causal manner, conditioning each new frame or segment on the historical context of previously generated content [chen2024diffusion, yin2025slow, huang2025self, liu2025rolling, yang2025longlive, chen2025skyreels, ji2025memflow, teng2025magi, NOVA, resampleforcing]. While this formulation naturally supports streaming generation with low initial latency, it is inherently susceptible to error accumulation, where minor deviations propagate and amplify across steps, leading to severe temporal drift and degraded coherence in long videos. To mitigate this issue, recent works introduce planning methods [zhang2025framepack, xiang2025macro] or explicit memory mechanisms [sun2025worldplay, chen2025teleworld, contex, mixtureof, memoryforcing, hyworld2025], strategies that typically necessitate complex architectural modifications and extensive re-training.
Test-time Image/Video Generation. Test-time generation methods [wang2025test, sun2020test, liang2025comprehensive] aim to enhance the performance of pre-trained models directly during the inference phase. Test-time scaling improves quality by iteratively searching over multiple candidates, as seen in Video-T1 [liu2025video] and EvoSearch [he2025scaling], yet this comes at the price of prohibitive computational cost. Similarly, Test-time optimization refines generation via auxiliary parameter updates that necessitate instance-specific training, such as HyperNoise [eyring2025noise], AutoRefiner [yu2025autorefiner], and SLOWFAST-VGEN [hong2025slowfastvgen]. In contrast, our approach distinguishes itself from both paradigms by being fully training-free, avoiding both the overhead of candidate search and the complexity of parameter optimization.
@@ 3.1 Background: Few-step Distilled Sampling
Under this formulation, conditional distribution $p_{\theta}(x_{t}\mid S_{t})$ is no longer realized as a single deterministic mapping, but through a stochastic diffusion sampling trajectory defined over a sparse set of diffusion steps $\{T_{0}=0,T_{1},\dots,T_{K}=T_{\max}\}$. Specifically, distilled video generation begins from Gaussian noise $x_{t}^{T_{\max}}\sim\mathcal{N}(0,I)$ and evolves progressively along this trajectory via a sequence of denoise–renoise transitions. Specifically, at each denoising step $T_{j}$, the generation process starts from a noisy latent state and applies the denoising network to produce an estimate of the underlying clean latent representation,
@@
Figure 4: Comparison of two toy test-time optimization variants based on LoRA fine-tuning, illustrating distinct failure modes.
3.2 Toy Experiment: Applying Test-time Optimization to Long Video Generation
Existing TTO methods improve generation quality by aligning the model distribution with a predefined reward function. Given a pre-trained generative model with output distribution $p^{\mathrm{base}}$, TTO typically defines a reward-weighted target distribution
@@ 3.2 Toy Experiment: Applying Test-time Optimization to Long Video Generation
To assess this idea, we conduct two toy experiments using direct LoRA fine-tuning [LoRA] at test time, following HyperNoise [eyring2025noise] and AutoRefiner [yu2025autorefiner]. Both variants use the same backbone and identical LoRA adapters, and differ only in their optimization objectives. The first variant fine-tunes LoRA with a standard denoising reconstruction loss on early frames across noise levels. The second variant replaces pixel-level reconstruction with a semantic consistency objective, enforcing similarity to early frames in pretrained feature spaces [CLIP, dinov2].
@@ 4.1 Correctability along the Stochastic Sampling Path
Motivated by planning-based video generation models [zhang2025framepack, xiang2025macro], which relax strictly unidirectional prediction via cross-frame context, we consider applying test-time intervention at a single sampling step after structural stabilization. Specifically, at a designated step $j^{\star}$, we restrict the visible context state $S_{t}$ to include only the earliest frame, forcing the model to rely exclusively on the initial frame for subsequent appearance refinement and texture generation. This single-point correction process can be formalized as:
@@
4.2 Pathwise Test-time Correction
@@
Algorithm 1 Pathwise Test-time Correction
[…]
@@
3: Sample $x_{t}^{T_{J}}\sim\mathcal{N}(0,I)$ # Initial gaussian noise
4:for $j=J$ down to 11 do
5:  $x_{t,0}^{T_{j}}\leftarrow{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}G_{\theta}}\!\left(x_{t}^{T_{j}};\,{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}S_{t}},\,j\right)$ # Initial prediction with $S_{t}$
@@
9:   $x_{t}^{T_{j-1},c}\leftarrow{\color[rgb]{0.21,0.12,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.21,0.12,1}\pgfsys@color@cmyk@stroke{0.79}{0.88}{0}{0}\pgfsys@color@cmyk@fill{0.79}{0.88}{0}{0}\Psi}\!\left(x_{t,0}^{T_{j}},\,\epsilon_{t}^{T_{j-1}},\,T_{j-1}\right)$
10:   $x_{t,0}^{T_{j-1},c}\leftarrow{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}G_{\theta}}\!\left(x_{t}^{T_{j-1},c};\,{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}S_{0}},\,j-1\right)$ # Correct trajectory using $S_{0}$
@@
13:   $x_{t}^{T_{j-1}}\leftarrow{\color[rgb]{0.21,0.12,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.21,0.12,1}\pgfsys@color@cmyk@stroke{0.79}{0.88}{0}{0}\pgfsys@color@cmyk@fill{0.79}{0.88}{0}{0}\Psi}\!\left(x_{t,0}^{T_{j-1},c},\,\tilde{\epsilon}_{t}^{T_{j-1}},\,T_{j-1}\right)$ # Inject new gaussian noise
14:   $x_{t,0}^{T_{j-1}}\leftarrow{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}G_{\theta}}\!\left(x_{t}^{T_{j-1}};\,{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}S_{t}},\,j-1\right)$ # Finalize step with $S_{t}$
@@
16:   $x_{t}^{T_{j-1}}\leftarrow{\color[rgb]{0.21,0.12,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.21,0.12,1}\pgfsys@color@cmyk@stroke{0.79}{0.88}{0}{0}\pgfsys@color@cmyk@fill{0.79}{0.88}{0}{0}\Psi}\!\left(x_{t,0}^{T_{j}},\,\epsilon_{t}^{T_{j-1}},\,T_{j-1}\right)$
17:   $x_{t,0}^{T_{j-1}}\leftarrow{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}G_{\theta}}\!\left(x_{t}^{T_{j-1}};\,{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}S_{t}},\,j-1\right)$
@@
20:return $x_{t,0}^{T_{0}}$
Table 1: Comprehensive comparison with SOTA methods on prompt-conditioned 30-second video generation (Speed & VBench).
Method Training Free Speed VBench Metrics $\uparrow$
Total fps Subject Consistency Background Consistency Dynamic Degree Motion Smoothness Imaging Quality Aesthetic Quality
Rolling Forcing 15.38 95.8 95.1 35.9 98.9 72.5 63.6
LongLive - 95.5 95.4 44.5 98.8 71.7 65.0
CausVid 15.79 91.2 91.4 50.8 98.1 70.2 63.5
CV + Ours 10.53 93.2 93.3 69.5 97.6 70.1 63.5
Self-Forcing 15.79 92.5 93.2 62.5 98.0 72.5 63.4
SF + Ours 10.53 94.0 94.2 60.2 98.3 72.7 63.8
Table 2: Comprehensive comparison with SOTA methods on prompt-conditioned 30-second video generation (Color-shift & JEPA). We report Color-shift metrics (L1, Correlation) and JEPA consistency (Variance, Difference).
Method Training Free Color-shift JEPA Consistency
L1 $\downarrow$ Correlation $\uparrow$ Variance $\downarrow$ Difference $\downarrow$
Rolling Forcing 0.436 0.858 0.0162 0.201
LongLive 0.701 0.724 0.0151 0.101
CausVid 1.047 0.451 0.0199 0.313
CV + Ours 0.607 0.778 0.0157 0.164
Self-Forcing 1.028 0.479 0.0145 0.191
SF + Ours 0.644 0.710 0.0108 0.170
Table 3: Comprehensive comparison with test-time scaling methods, including Best-of-N (BoN) and Search-over-Path (SoP), on prompt-conditioned 30-second video generation. We report Throughput (fps) and VBench metrics.
Method Train. Free Speed VBench Metrics $\uparrow$
Total fps Subject Consistency Background Consistency Dynamic Degree Motion Smoothness Imaging Quality Aesthetic Quality